OpenStudy (anonymous):

Help! put the equation in slope-intercept for by solving for y. 4y = 5x

OpenStudy (anonymous):

assuming you mean slope intersect form as in y=m*x+b, we know that b=0, so all we need to do is divide by 4 on either side, there by giving us y=(5/4)x

OpenStudy (anonymous):

I DID THAT!!! but I was not sure. thank you very much!

OpenStudy (anonymous):

Yep, anytime :)